# Partial confinement in a quantum-link simulator

### v1

@@ I Introduction
Confinement is a fundamental property prominently observed in quantum chromodynamics (QCD), where the inter-quark potential increases with their distance [1, 2, 3]. This prevents the existence of isolated quarks due to energetic instability; instead, they prefer to bind together into hadrons, either as mesons (quark-antiquark pairs) or baryons (triplets of quarks). Although the concept originated in QCD, analogous phenomena can also manifest in strongly coupled charges in quantum electrodynamics (QED) [4, 5], i.e., the charge confinement. For instance, in (3+1)D QED, increasing the coupling strength $g$ facilitates a transition from the deconfined Coulomb phase to the confining phase at finite temperature [6]. Particularly in (1+1)D QED, also known as the Schwinger model, the absence of a magnetic field renders the system to be always strongly coupled. Consequently, apart from certain exceptional cases, the confining phase becomes quite prevalent even at zero temperature [7, 8]. Confinement and deconfinement phenomena also appear in emergent gauge theories in the strongly-correlated quantum matter. For instance, transition between valence bond solid to spin liquid phase can be understood in a picture of confinement-deconfinement transition of spinons [9, 10]. However, large scale numerical investigation of the real time dynamics of confinement or deconfinement on classical computers is challenging.
@@
Recently, much efforts have been made to overcome this barrier via the approach of quantum simulation, which relies on systems with discrete degrees of freedom, such as cold atoms in optical lattices [11, 12, 13, 14, 15] or trapped ions [16, 17]. The simulation implements the lattice gauge theory through a Hamiltonian formalism called the quantum link model (QLM) [18, 19, 20, 21, 22], with matter particles and gauge spins being respectively placed on lattice sites and links connecting neighboring sites. The realization of QLM is considered as a powerful approach for exploring strongly coupled QED [18], as strong coupling renders perturbative field theory ineffective, and hence quantum simulation can essentially circumvent the issues encountered by classical simulations, such as the sign problem in quantum Monte Carlo [23]. In these quantum-link simulators, both confinement and deconfinement have been extensively studied, encompassing theoretical [24, 25, 26, 27, 28, 29, 30, 31] and experimental [32, 15] contexts. Particularly for the spin-1/2 QLM, the confinement-deconfinement phase transition has been experimentally observed in non-equilibrium dynamics by tuning the topological angle [15].
@@
In this paper, we delve into an intermediate phenomenon between confinement and deconfinement, called partial confinement, which has hitherto remained unexplored within the context of quantum simulation. It refers to the situation where the confining or deconfining status between charges depends on their relative positions. Our study draws inspiration from the seminal work of S. Coleman in the 1976 [5], which studied half-asymptotic particles in the continuous Schwinger model. Here, we demonstrate that the one-dimensional spin-1 QLM can serve as an excellent platform for observing partial confinement: it retains the essential physics while being simple enough to be realized within the scope of current experimental capabilities. Taking the spin-1 QLM as a background, we introduce the basic concept of partial confinement and discuss the associated emergent physics, both in equilibrium and non-equilibrium dynamics.
@@ II Spin-1 Quantum Link Model
where ${\psi}_{j}$ denotes the local matter fields of fermions, and ${S}_{j}^{z,\pm}$ are the spin-1 Pauli operators representing the gauge spins living on the link between two neighboring sites $j$ and $j+1$. The number of sites $N$ must be even, allowing for the division of fermions into positrons and electrons, yielding the particle-antiparticle picture: for $j\in\text{even}$, $|0\rangle$ and $|1\rangle={\psi}_{j\in\text{even}}^{\dagger}|0\rangle$ respectively represent the unoccupied and occupied states of electrons, denoted by empty circles and red disks in Fig. 1; whereas for $j\in\text{odd}$, $|1\rangle$ represents the occupation of positrons, illustrated by blue disks.
@@
In some literature [24, 37, 38, 39], the QLM [Eq. (1)] is presented in an alternative form within the particle picture, described by the Hamiltonian
@@
which relates to the particle-antiparticle Hamiltonian $H$ [Eq. (1)] through a particle-hole transformation on odd sites, i.e. $\psi_{j\in\text{odd}}\mapsto\psi^{\dagger}_{j\in\text{odd}}$, as well as a transformation on even gauge spin $S^{+}_{j\in\text{even}}\mapsto-S^{-}_{j\in\text{even}}$ and $S^{z}_{j\in\text{even}}\mapsto-S^{z}_{j\in\text{even}}$. In this framework, $\psi^{\dagger}_{j}\psi_{j}$ at odd sites represents the occupation below the Dirac sea, thereby exhibiting a negative energy (mass) $-m$. The creation of a hole $\psi_{j}$ in the particle picture is equivalent to the creation of a positron $\psi_{j}^{\dagger}$ in the particle-antiparticle picture. Note that both Hamiltonians, $H$ and $\tilde{H}$, are mathematically equivalent for calculation purposes. Therefore, we proceed with our following analysis using the particle-antiparticle picture.
@@ III Partial Confinement in Equilibrium
The confining effects can be clearly demonstrated by the properties of equilibrium states in the physical sector $\mathbf{q}=\mathbf{0}$. We first focus on the simplest case of $J=0$, where the matter and gauge fields are decoupled, thereby $\psi^{\dagger}_{j}\psi_{j}$ and $S_{j}^{z}$ being conserved. We insert a pair of testing electron and positron into the vacuum, separated by a distance $d$, as schematically shown in Fig. 2(a). $d$ can be positive or negative, with $d>0$ indicating the positron is to the left of the electron, and vice versa. The case of $d=0$ is excluded by the Pauli exclusion principle. According to Gauss’s Law [Eq. (3)], the system is in the string state
@@
Figure 3: Behaviors of the states’ energy $\mathcal{E}$ and string tension $\rho$ in equilibrium with fixed $N=16$, $g=1$ and $m=6g$. Panels (a)-(d) correspond to the integrable case $J=0$: (a) displays the dependence of the string state energy $\mathcal{E}_{\text{str}}$ and the meson state energy $\mathcal{E}_{\text{meson}}$ on $d$ for $\theta=0$; (b) depicts the behavior of $\mathcal{E}_{\text{str}}$ and $\mathcal{E}_{\text{meson}}$ for $\theta=\pm\pi$; (c) shows $\mathcal{E}_{\text{str}}$ for different values of $\theta$; (d) presents the string tension $\rho$ as a function of the topological angle $\theta$ for fixed $d=\pm 12$. Panels (e)-(f) correspond to the non-integrable case $J=1$: (e), analogous to panel (b), illustrates the string-like eigenstate energy $\mathcal{E}_{\text{str}}$ and the meson-like eigenstate energy $\mathcal{E}_{\text{meson}}$ for $\theta=\pm\pi$; (f), analogous to panel (d), shows the string tension as a function of the topological angle $\theta$ for fixed $d=\pm 12$.
@@
The above results for $J=0$ will not be qualitatively altered when we turn on the matter-gauge interaction $J$. When $J$ is finite, the system lacks integrability, causing a resort to numerical calculations. By setting $J=g=1$ and $m=6g$, we numerically calculate the energy spectrum of the system. Although quantum fluctuations render $s_{j}^{z}$ and $n_{j}$ no longer good quantum numbers, we can still identify low-energy string-like and meson-like states, with the averaged local observables, such as $\langle S_{j}^{z}\rangle$ and $\langle\psi_{j}^{\dagger}\psi_{j}\rangle$, resembling the configurations of string and meson states shown in Fig. 3(b). Intuitively, the string remains but is thickened by quantum fluctuations. In Fig. 3(e), we present the energy variance of these two eigenstates $\mathcal{E}$ on $d$ for various $\theta$; Additionally, Fig. 3(f) shows the corresponding string tension $\rho$ as a function of $\theta$ for a fixed $d=\pm 12$. A comparison between panels (b) and (e), as well as (d) and (f), clearly demonstrates that the main physical results, such as partial confinement and string breaking, are qualitatively preserved for a non-vanishing $J$.
It may be also necessary to elucidate the differences between the spin-1 QLM discussed here and the spin-1/2 QLM, with the latter, has been extensively studied both theoretically and experimentally [12, 13, 14, 26, 27, 15, 24, 25, 40, 41]. The Hamiltonian of the spin-1/2 QLM has the same form as Eq. (1), but with the spin operators $S_{j}$ being spin-1/2 Pauli operators (up to a constant factor). In this case, even without an external electric field, the string state is no longer well-defined, as there always exist electric strings between the charges (inner string) and outside the charges (outer string). Commonly, the spin-1/2 QLM with the outer string pointing to the right (left) is considered to have an inherent topological angle $\theta=\pi$ ($\theta=-\pi$) [25, 26, 27, 15]. In contrast to the spin-1 case, the charge conjugation $\mathcal{C}$ now would simultaneously change the order of the charges and the sign of $\theta$, rendering the total energy $\mathcal{E}$ irrelevant to the sign of $d$. In this sense, the spin-1 QLM may serve as a better platform for the study of partial confinement, as it allows for independently changing the charge ordering and $\theta$.
@@ IV Partial Confinement in dynamics
After discussing equilibrium physics in depth, we now turn to non-equilibrium dynamics. Our objective is to explore whether the physics of partial confinement can be signified by the quantum dynamics out of equilibrium. To this end, we initialize the state $|\psi_{0}\rangle$ to be a string state with the positron and electron residing on the two edges of the chain with $d>0$, i.e.,
@@
Notably, this state is an eigenstate of the Hamiltonian $H$ when $J=0$. We then allow the system to evolve under the government of $H$ with $J=g=1$. In our numerical simulations, $N=16$, $m=2g$ are fixed. We primarily focus on the three cases of $\theta=\{-\pi,0,\pi\}$. According to the previous discussions, for such a string configuration $|\psi_{0}\rangle$ with the positron situated to the left of the electron, both $\theta=0$ and $\theta=-\pi$ are confining, while $\theta=\pi$ is deconfining. Since flipping the sign of $\theta$ is equivalent to changing the sign of charge ordering $d$, as discussed in Sec. III, comparing the dynamics at ±$\pm\theta$ for a fixed string state can directly signify the partial confinement.
We begin by examining the expectation values of fermion occupations $\langle n_{j}\rangle=\langle\psi_{j}^{\dagger}\psi_{j}\rangle$ within the time frame $t\leq 30g^{-1}$, as shown in Fig. 4 with panels (a), (b), and (c) corresponding to the cases of $\theta=-\pi$, 00, and $\pi$. In the figures, the occupations of electrons and positrons are labeled by red and blue color bars, respectively. A prominent feature is that, for the first two confining cases (i.e., $\theta=-\pi$ and 00), the edge charges are ’locked’ at the boundaries with almost no movement. In the bulk of the chain, $\langle n_{j}\rangle$ overall exhibits a periodic oscillation, which is indicative of the Schwinger mechanism, as labeled by the white boxes: electron-positron pairs are spontaneously created from the vacuum and then rapidly annihilated with each other. The confinement effect is also evidenced by the small lifetime of the emerged particles (anti-particles) and the fact that they cannot propagate to a wider range on the chain.
@@
Figure 4: (a)-(c) Dynamics of fermion occupations $\langle n_{j}\rangle$ for $\theta=-\pi$ [panel (a)], $\theta=0$ [panel (b)], and $\theta=\pi$ [panel (c)], where the occupation status of the electron and positron are labeled by the red and blue color bars, respectively, the white boxes indicate the Schwinger mechanism. (d) Top 10 projection probabilities $|c_{n}|^{2}$ of the initial state $|\psi_{0}\rangle$ onto the eigenstates $|n\rangle$ of $H$ under different $\theta$ value. In the calculation, we take $N=16$, $m=2g$, and $J=g=1$.
@@
On the other hand, the case with $\theta=\pi$ [Fig. 4(c)] exhibits a strikingly different behavior. The two edge charges move towards each other until they meet at the center of the chain at about $t\approx 12g^{-1}$; after that, they reverse their directions and retreat to the boundaries. This exemplifies partial confinement effectively. At the outset, due to $\theta=\pi$, the electron and positron are deconfined, which permits their free movement toward the chain’s center. As they exchange their positions at about $j=N/2$, the charges become confined. The confinement prevents their further propagation, compelling them to reverse direction. Additionally, it can be also notable from the figure that the particle occupation $\langle n_{j}\rangle$ at the boundaries is significantly decreased compared to the initial state after one round trip. Unlike the Schwinger oscillations in the former two cases, the back-and-forth motion of particles can now only sustain for a few cycles and lacks robust periodicity.
@@
Figure 5: (a) Dynamics of the bipartite entanglement entropy $S$ for various cases of $\theta$. For the confining cases, $S$ exhibits a periodic oscillations and grows slowly due to Schwinger oscillation. On contrary, for the deconfining case, $S$ presents a rapid increase. (b) Dynamics of the density-density correlator $\langle n_{1}n_{N}\rangle_{\text{con}}$ between the two boundary fermions, for the confining cases, only weak correlation persisted. Oppositely, for deconfining case, correlation can grow rapidly. The parameters take $N=16$, $m=2J$, and $J=g=1$.
We additionally calculate the dynamics of the connected density correlation, defined by

### v2

@@ I Introduction
Confinement is a fundamental property prominently observed in quantum chromodynamics (QCD), where the inter-quark potential increases with their distance [1, 2, 3]. This prevents the existence of isolated quarks due to energetic instability; instead, they prefer to bind together into hadrons, either as mesons (quark-antiquark pairs) or baryons (triplets of quarks). Although the concept originated in QCD, analogous phenomena can also manifest in strongly coupled charges in quantum electrodynamics (QED) [4, 5], i.e., the charge confinement. Dimensional analysis indicates that the dimensionality of the coupling constant is determined by the dimensions of the system. Specifically for (3+1)D, the coupling constant is dimensionless, leading to the deconfined Coulomb potential $\sim 1/r$, where $r$ is the distance between two charges. The deconfinement-confinement phase transition can occur by tuning the coupling strength and the temperature [6]. However, for (1+1)D QED, also known as the Schwinger model, the dimensionality of the coupling constant scales linearly with $r$. Consequently, apart from certain exceptional cases, the confining phase becomes quite prevalent. Furthermore, confinement and deconfinement phenomena also appear in emergent gauge theories from strongly-correlated electrons and recent developed Rydberg atomic arrays [7, 8, 9, 10, 11]. For instance, transition between valence bond solid to spin liquid phase can be understood in a picture of confinement-deconfinement transition of spinons [12, 13]. However, large scale numerical investigation of the real time dynamics of confinement or deconfinement on classical computers is challenging.
Recently, much effort has been made to overcome this barrier through quantum simulation [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] which leverages systems with discrete degrees of freedom. This includes analog simulations using optical lattices [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46] or trapped ions [47, 48, 49, 50], and digital simulations realized on various quantum computing platforms [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63].
@@
The quantum link model (QLM) [64] serves as one of the most commonly used approaches to simulate lattice gauge theories, which are based on the Hamiltonian formalism with space discretized while time remains continuous. In QLMs, matter particles are placed on lattice sites, while gauge spins with a finite local Hilbert space are located on the links connecting neighboring sites. The realization of QLM is considered a powerful approach for exploring strongly coupled QED, as strong coupling renders perturbative field theory ineffective, and hence quantum simulation can essentially circumvent the issues encountered by classical simulations, such as the sign problem in quantum Monte Carlo [65]. In these quantum-link simulators, both confinement and deconfinement have been extensively studied, encompassing theoretical [44, 66, 67, 68, 69, 70, 71, 72, 73, 45] and experimental [74, 30] contexts. Particularly for the spin-1/2 QLM, the confinement-deconfinement transition has been experimentally signified in dynamics through tuning the topological angle [30].
@@
In this paper, we delve into an intermediate phenomenon between confinement and deconfinement in (1+1)D QED, called partial confinement, which has hitherto remained unexplored within the context of quantum simulation. It refers to the situation where the confining or deconfining status between charges depends on their relative positions. Our study draws inspiration from the seminal work of S. Coleman in the 1976 [5], which studied half-asymptotic particles in the continuum Schwinger model. Here, we demonstrate that the 1D spin-1 QLM can serve as an excellent platform for observing partial confinement: it retains the essential physics while being simple enough to be realized within the scope of current experimental capabilities. Taking the spin-1 QLM as a background, we introduce the basic concept of partial confinement and discuss the associated emergent physics, both in equilibrium and non-equilibrium dynamics.
It is worth clarifying that the terminology of partial (de)confinement has already been introduced in QCD [75, 76, 77]. Therein, partial confinement refers to a phase where color degrees of freedom split into confined and deconfined sectors, with a subgroup of the SU(N) gauge group becoming deconfined while the remainder stays confined. This is typically characterized by a non-uniform distribution of Polyakov loop phases and the N-dependence of thermodynamic quantities. As such, this notion of partial confinement carries a different physical meaning compared to ours defined above.
@@ II Spin-1 Quantum Link Model
where ${\psi}_{j}$ denotes the local matter fields of fermions, and ${S}_{j}^{z,\pm}$ are the spin-1 Pauli operators representing the gauge spins living on the link between two neighboring sites $j$ and $j+1$. The number of sites $N$ must be even, allowing for the division of fermions into electrons and positrons, yielding the particle-antiparticle picture: for $j\in\text{odd}$, the unoccupied and occupied status of electrons are respectively denoted by empty circles and blue disks in Fig. 1; whereas for $j\in\text{even}$, the corresponding occupation status of positrons are illustrated by empty circles and red disks.
@@
In some literature [44, 82, 48, 83], the QLM [Eq. (1)] is presented in an alternative form within the particle picture, described by the Hamiltonian with staggered mass [84, 85]
@@
which relates to the particle-antiparticle Hamiltonian $H$ [Eq. (1)] through a particle-hole transformation on odd sites, i.e. $\psi_{j\in\text{odd}}\mapsto\psi^{\dagger}_{j\in\text{odd}}$, as well as a transformation on even gauge spin $S^{+}_{j\in\text{even}}\mapsto-S^{-}_{j\in\text{even}}$ and $S^{z}_{j\in\text{even}}\mapsto-S^{z}_{j\in\text{even}}$. In this framework, $\psi^{\dagger}_{j}\psi_{j}$ at odd sites represents the occupation below the Dirac sea, thereby exhibiting a negative energy (mass) $-m$. The creation of a hole $\psi_{j}$ in the particle picture is equivalent to the creation of an electron $\psi_{j}^{\dagger}$ in the particle-antiparticle picture. Note that both Hamiltonians, $H$ and $\tilde{H}$, are mathematically equivalent for calculation purposes. Therefore, we proceed with our following analysis using the particle-antiparticle picture.
@@ III Partial Confinement in Equilibrium
The confining effects can be clearly demonstrated by the properties of equilibrium states in the physical sector $\mathbf{q}=\mathbf{0}$. We first focus on the simplest case of $J=0$, where the matter and gauge fields are decoupled, thereby $\psi^{\dagger}_{j}\psi_{j}$ and $S_{j}^{z}$ being conserved. We insert a pair of test electron and positron into the vacuum, separated by a distance $d$, as schematically shown in Fig. 2(a). $d$ can be positive or negative, with $d>0$ indicating the electron is to the left of the positron, and vice versa. The case of $d=0$ is excluded by the Pauli exclusion principle. According to Gauss’s Law [Eq. (3)], the system is in the string state
@@
Figure 3: Equilibrium states’ energy $\mathcal{E}$ and string tension $\rho$ in the case of $J=0$, with $N=16$, $g=1$ and $m=6g$ being fixed. Panel (a) displays the dependence of the string state energy $\mathcal{E}_{\text{str}}$ and the meson state energy $\mathcal{E}_{\text{meson}}$ on $d$ for $\theta=0$; (b) depicts the behavior of $\mathcal{E}_{\text{str}}$ and $\mathcal{E}_{\text{meson}}$ for $\theta=\pm\pi$; (c) shows $\mathcal{E}_{\text{str}}$ for different values of $\theta$; (d) presents the string tension $\rho$ as a function of the topological angle $\theta$ for fixed $d=\pm 12$.
@@
Figure 4: Equilibrium states’ energy $\mathcal{E}$ and string tension $\rho$ in the case of $J=1$, with all other parameters being the same with Fig. 3. Panel (a) illustrates the string-like eigenstate energy $\mathcal{E}_{\text{str}}$ and the meson-like eigenstate energy $\mathcal{E}_{\text{meson}}$ for $\theta=\pm\pi$. Panel (b) shows the string tension as a function of the topological angle $\theta$ for fixed $d=\pm 12$.
The above results for $J=0$ will not be qualitatively altered when we turn on the matter-gauge interaction $J$. When $J$ is finite, the system lacks integrability, causing a resort to numerical calculations. By setting $J=g=1$ and $m=6g$, we numerically calculate the energy spectrum of the system. Although quantum fluctuations render $s_{j}^{z}$ and $n_{j}$ no longer good quantum numbers, we can still identify low-energy string-like and meson-like states, with the averaged local observables, such as $\langle S_{j}^{z}\rangle$ and $\langle\psi_{j}^{\dagger}\psi_{j}\rangle$, resembling the configurations of string and meson states shown in Fig. 3(b). Intuitively, the string remains but is thickened by quantum fluctuations. In Fig. 4(a), we present the energy variance of these two eigenstates $\mathcal{E}$ on $d$ for various $\theta$; Additionally, Fig. 4(b) shows the corresponding string tension $\rho$ as a function of $\theta$ for a fixed $d=\pm 12$. A comparison between panels Fig. 3(b) and Fig. 4(a), as well as Fig. 3(d) and Fig. 4(b), clearly demonstrates that the main physical results, such as partial confinement and string breaking, are qualitatively preserved for a non-vanishing $J$.
It may also be necessary to elucidate the differences between the spin-1 QLM discussed here and the spin-1/2 QLM which has been extensively studied both theoretically and experimentally [27, 28, 29, 66, 67, 30, 44, 10, 86, 87]. The Hamiltonian of the spin-1/2 QLM has the same form as Eq. (1), but with the spin operators $S_{j}$ being spin-1/2 Pauli operators (up to a constant factor). In this case, even without an external electric field, the string state is no longer well-defined, as there always exist electric strings between the charges (inner string) and outside the charges (outer string). Commonly, the spin-1/2 QLM with the outer string pointing to the right (left) is considered to have an inherent topological angle $\theta=\pi$ ($\theta=-\pi$) [10, 66, 67, 30]. In contrast to the spin-1 case, the charge conjugation $\mathcal{C}$ now would simultaneously change the order of the charges and the sign of $\theta$, rendering the total energy $\mathcal{E}$ irrelevant to the sign of $d$. In this sense, the spin-1 QLM may serve as a better platform for the study of partial confinement, as it allows for independently changing the charge ordering and $\theta$.
@@ IV Partial Confinement in dynamics
IV.1 String-state dynamics
After discussing equilibrium physics in depth, we now turn to non-equilibrium dynamics. Our objective is to explore whether the physics of partial confinement can be signified by the quantum dynamics out of equilibrium. To this end, we first consider the initial state $|\psi_{0}\rangle$ to be a string state with the electron and positron residing on the two edges of the chain with $d>0$, i.e.,
@@ IV.1 String-state dynamics
Notably, this state is an eigenstate of the Hamiltonian $H$ when $J=0$. We then allow the system to evolve under the government of $H$ with $J=g=1$. In our numerical simulations, $N=16$, $m=2g$ are fixed. We primarily focus on the three cases of $\theta=\{-\pi,0,\pi\}$. According to the previous discussions, for such a string configuration $|\psi_{0}\rangle$ with the electron situated to the left of the positron, both $\theta=0$ and $\theta=-\pi$ are confining, while $\theta=\pi$ is deconfining. Since flipping the sign of $\theta$ is equivalent to changing the sign of charge ordering $d$, as discussed in Sec. III, comparing the dynamics at ±$\pm\theta$ for a fixed string state can directly signify the partial confinement.
@@
Figure 5: Density dynamics of the string state. (a)-(c) Dynamics of fermion occupations $\langle n_{j}\rangle$ for $\theta=-\pi$ [panel (a)], $\theta=0$ [panel (b)], and $\theta=\pi$ [panel (c)], where the occupation status of the positrons and electrons are labeled by the red and blue color bars, respectively, the white boxes indicate the Schwinger mechanism. (d) Top 10 projection probabilities $|c_{n}|^{2}$ of the initial state $|\psi_{0}\rangle$ onto the eigenstates $|n\rangle$ of $H$ under different $\theta$ value. In the calculation, we take $N=16$, $m=2g$, and $J=g=1$.
We begin by examining the expectation values of fermion occupations $\langle n_{j}\rangle=\langle\psi_{j}^{\dagger}\psi_{j}\rangle$ within the time frame $t\leq 30g^{-1}$, as shown in Fig. 5 with panels (a), (b), and (c) corresponding to the cases of $\theta=-\pi$, 00, and $\pi$. In the figures, the occupations of positrons and electrons are labeled by red and blue color bars, respectively. A prominent feature is that, for the first two confining cases (i.e., $\theta=-\pi$ and 00), the edge charges are ’locked’ at the boundaries with almost no movement. In the bulk of the chain, $\langle n_{j}\rangle$ overall exhibits a periodic oscillation, which is indicative of the Schwinger mechanism, as labeled by the white boxes: electron-positron pairs are spontaneously created from the vacuum and then rapidly annihilated with each other. The confinement effect is also evidenced by the small lifetime of the emerged particles (anti-particles) and the fact that they cannot propagate to a wider range on the chain.
@@
On the other hand, the case with $\theta=\pi$ [Fig. 5(c)] exhibits a strikingly different behavior. The two edge charges move towards each other until they meet at the center of the chain at about $t\approx 12g^{-1}$; after that, they reverse their directions and retreat to the boundaries. Unlike the traditional scattering process for free fermions where transmitted waves continue to propagate forward after the scattering event, here we do not observe a clear signal of transmitted wave propagation. Additionally, it can be also notable from the figure that the particle occupation $\langle n_{j}\rangle$ at the boundaries is significantly decreased compared to the initial state after one round trip. Unlike the Schwinger oscillations in the former two cases, the back-and-forth motion of particles can now only sustain for a few cycles and lacks robust periodicity.
@@
Figure 6: Dynamics of the entanglement and the correlator for the string state. (a) Dynamics of the bipartite entanglement entropy $S$ for various cases of $\theta$. For the confining cases, $S$ exhibits a periodic oscillations and grows slowly due to Schwinger oscillation. On contrary, for the deconfining case, $S$ presents a rapid increase. (b) Dynamics of the density-density correlator $\langle n_{1}n_{N}\rangle_{\text{con}}$ between the two boundary fermions, for the confining cases, only weak correlation persisted. Oppositely, for deconfining case, correlation can grow rapidly. The parameters take $N=16$, $m=2J$, and $J=g=1$.
@@
We also calculate the dynamics of the connected density correlation, defined by
@@
IV.2 Meson-state dynamics
Up to now, our discussion has focused on the quench dynamics of the string state. Here, we additionally consider a scenario where the initial state exhibits a single meson excitation with $d>0$ at the center of the ground state, i.e.,
$|\psi_{0}\rangle=|\psi_{\text{meson}}\rangle=\psi^{\dagger}_{\frac{N}{2}}S^{-}_{\frac{N}{2}}\psi^{\dagger}_{\frac{N}{2}+1}|\psi_{g}\rangle,$ (14)
where the $|\psi_{g}\rangle$ is the ground state of the Hamiltonian $H$ [Eq. (1)] for $m=4g$ and $J=g=1$. The additional gauge-spin flip $S^{-}_{\frac{N}{2}}$ ensures that $|\psi_{0}\rangle$ remains within the physical gauge sector. Since $d>0$, the particle-antiparticle pair is confined for $\theta=\{-\pi,0\}$ and deconfined for $\theta=\pi$. We simulate the dynamics of $|\psi_{0}\rangle$, with the results of the fermion occupations $\langle n_{j}\rangle$ being presented in Fig. 7. Again, the panels (a)-(c) correspond to the cases of $\theta=\{-\pi,0,\pi\}$ respectively.
Figure 7: Density dynamics of the meson state in the cases of $\theta=-\pi$ [panel (a)], $\theta=0$ [panel (b)], and $\theta=\pi$ [panel (c)], where the occupation status of the positrons and electrons are labeled by the red and blue color bars, respectively. In the calculation, we take $N=14,m=4g$ and $J=g=1$.
One can observe that $\langle n_{j}\rangle$ exhibits distinct behaviors for different values of $\theta$. Specifically, in the case of strong confinement at $\theta=-\pi$, the meson state remains stuck in the center of the chain without movement. For the relatively weaker confining case with $\theta=0$, the meson simultaneously moves towards both ends of the chain, ensuring conservation of momentum. Thus, for these two confining cases, the positron and electron are bound together. In contrast, for the deconfining case of $\theta=\pi$, the meson dissociates into isolated positron and electron, which then independently move away from each other.
The distinction between the confining and deconfining cases can also manifest in the correlation function [66]
$\mathcal{G}(r,t)=\sum_{j=1}^{N}\langle\psi_{0}(t)|(n_{j}-\bar{n}_{j})(n_{j+r}-\bar{n}_{j+r})|\psi_{0}(t)\rangle,$ (15)
where $n_{j}=\psi_{j}^{\dagger}\psi_{j}$ and $\bar{n}_{j}=\langle\psi_{g}|n_{j}|\psi_{g}\rangle$. $\mathcal{G}(r,t)$ quantifies the spatial correlation of density fluctuations between any two arbitrary lattice sites separated by a distance $r$. The numerical results of $\mathcal{G}(r,t)$ for various $\theta$ are respectively shown in Fig. 8. Specifically, for the confining cases with $\theta=\{0,-\pi\}$, the correlation function $\mathcal{G}(r,t)$ remains localized and does not diffuse as $t$ increases. In contrast, for the deconfining case with $\theta=\pi$, the correlation function between the positron and electron can propagate at a considerable speed and eventually spread throughout the entire space.
Figure 8: Dynamics of correlation function for the meson state in the case of $\theta=-\pi$ [panel (a)], $\theta=0$ [panel (b)], and $\theta=\pi$ [panel (c)], with all the parameters being the same with Fig, 7.